ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: Tomáš Roubíček, Charles University, Prague Lebesgue spaces, typically involved in various optimization problems, can be homeomorphically and densely imbedded into convex locally compact envelopes. This enables to extend continuously various optimization problems and treat them like in a finite-dimensional case. The talk is based on the recently appeared book of the same title (W. de Gruyter, Berlin, 1997).